# GradMix: Gradient-based Selective Mixup for Robust Data Augmentation in Class-Incremental Learning

### v1

@@ C.5 More Details on Computational Complexity and Runtime Results of GradMix
Figure 11: Runtime results for solving the optimization problem in GradMix per epoch with respect to the number of tasks on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
In addition, we provide the average training time results over all tasks, both without and with GradMix, based on the experience replay method (ER) across all datasets as shown in Table 8. The computational overhead of GradMix depends on the last layer gradient computation, but our method computes gradients of only a small exemplar set for each class, not the whole training data. While GradMix incurs some additional computational cost relative to conventional training, we believe this is a worthwhile trade-off as it consistently leads to improved overall accuracy.
Table 8: Average training time results over all tasks, both without and with GradMix, based on the experience replay method (ER) on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
@@
C.6 More Details on Sensitivity of GradMix to the Mixup Parameter $\alpha$
Continuing from Sec. 6, we perform additional experiments to analyze the sensitivity of GradMix to the Mixup parameter $\alpha$ as shown in Table 9. As a result, the average accuracy of GradMix is not sensitive to $\alpha$. We believe that mixing the helpful pairs of classes that do not cause catastrophic forgetting might be beneficial to stabilize its performance regardless of $\alpha$. For Tiny ImageNet, a small value of $\alpha$ shows better accuracy results as in the Mixup paper [75], and our selective mixup algorithm struggles to overcome the data characteristics.
Table 9: Average accuracy results of GradMix when varying the $\alpha$ value to 0.4, 1.0, and 2.0 on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
@@ C.6 More Details on Sensitivity of GradMix to the Mixup Parameter $\alpha$
GradMix ($\alpha=0.4$) 0.919$\pm$0.005 0.795$\pm$0.006 0.657$\pm$0.010 0.548$\pm$0.006 0.405$\pm$0.006
GradMix ($\alpha=1.0$) 0.918$\pm$0.004 0.802$\pm$0.009 0.667$\pm$0.025 0.546$\pm$0.008 0.386$\pm$0.006
GradMix ($\alpha=2.0$) 0.917$\pm$0.001 0.809$\pm$0.007 0.660$\pm$0.020 0.543$\pm$0.009 0.372$\pm$0.005
C.7 More Details on Compatibility of GradMix to Advanced Experience Replay Methods
Continuing from Sec. 6.1, we perform additional experiments when using MIR [3], GSS [4], and DER [7] as basis experience replay methods for class-incremental learning, and the results are shown in Table 10–Table 12. As a result, GradMix still outperforms other data augmentation baselines when based on advanced experience replay methods. We believe that GradMix also can be extended to other experience replay methods including ASER [58] and SCR [47].
Table 10: Average accuracy results on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. We use MIR as a basis method of class-incremental learning and compare GradMix with three types of data augmentation baselines: (1) mixup-based methods: Mixup, Manifold Mixup, and CutMix; (2) imbalance-aware mixup methods: Remix and Balanced-MixUp; and (3) a policy-based method: RandAugment.
@@ C.7 More Details on Compatibility of GradMix to Advanced Experience Replay Methods
Experience Replay MIR 0.900$\pm$0.006 0.752$\pm$0.006 0.620$\pm$0.009 0.493$\pm$0.021 0.386$\pm$0.009
Mixup-based + Mixup 0.878$\pm$0.010 0.770$\pm$0.006 0.644$\pm$0.008 0.504$\pm$0.011 0.372$\pm$0.009
+ Manifold Mixup 0.900$\pm$0.007 0.776$\pm$0.005 0.619$\pm$0.005 0.522$\pm$0.009 0.388$\pm$0.008
+ CutMix 0.884$\pm$0.005 0.787$\pm$0.006 0.514$\pm$0.010 0.406$\pm$0.010 0.349$\pm$0.006
Imbalance-aware + Remix 0.911$\pm$0.006 0.789$\pm$0.009 0.638$\pm$0.009 0.525$\pm$0.009 0.393$\pm$0.007
+ Balanced-MixUp 0.887$\pm$0.006 0.771$\pm$0.008 0.627$\pm$0.018 0.515$\pm$0.013 0.380$\pm$0.008
Policy-based + RandAugment 0.770$\pm$0.007 0.648$\pm$0.010 0.636$\pm$0.005 0.485$\pm$0.008 0.366$\pm$0.010
Selective Mixup + GradMix (ours) 0.916$\pm$0.003 0.797$\pm$0.003 0.669$\pm$0.013 0.546$\pm$0.007 0.411$\pm$0.009
Table 11: Average accuracy results on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. We use GSS as a basis method of class-incremental learning, and other experimental settings are the same as in Table 10.
@@
Experience Replay GSS 0.911$\pm$0.005 0.772$\pm$0.007 0.627$\pm$0.009 0.512$\pm$0.013 0.397$\pm$0.008
Mixup-based + Mixup 0.884$\pm$0.008 0.777$\pm$0.010 0.653$\pm$0.010 0.532$\pm$0.011 0.387$\pm$0.007
+ Manifold Mixup 0.912$\pm$0.006 0.781$\pm$0.006 0.628$\pm$0.009 0.543$\pm$0.007 0.404$\pm$0.007
+ CutMix 0.892$\pm$0.007 0.793$\pm$0.006 0.523$\pm$0.011 0.417$\pm$0.011 0.348$\pm$0.006
Imbalance-aware + Remix 0.904$\pm$0.007 0.797$\pm$0.008 0.647$\pm$0.010 0.549$\pm$0.008 0.406$\pm$0.006
+ Balanced-MixUp 0.886$\pm$0.007 0.776$\pm$0.007 0.642$\pm$0.012 0.538$\pm$0.010 0.393$\pm$0.007
Policy-based + RandAugment 0.768$\pm$0.009 0.659$\pm$0.008 0.643$\pm$0.008 0.497$\pm$0.008 0.381$\pm$0.009
Selective Mixup + GradMix (ours) 0.927$\pm$0.005 0.808$\pm$0.007 0.674$\pm$0.011 0.563$\pm$0.007 0.418$\pm$0.006
Table 12: Average accuracy results on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. We use DER as a basis method of class-incremental learning, and other experimental settings are the same as in Table 10.
@@
Experience Replay DER 0.927$\pm$0.004 0.794$\pm$0.007 0.678$\pm$0.010 0.560$\pm$0.009 0.422$\pm$0.008
Mixup-based + Mixup 0.902$\pm$0.006 0.801$\pm$0.009 0.698$\pm$0.011 0.574$\pm$0.008 0.413$\pm$0.007
+ Manifold Mixup 0.927$\pm$0.005 0.810$\pm$0.005 0.687$\pm$0.008 0.580$\pm$0.006 0.429$\pm$0.007
+ CutMix 0.913$\pm$0.005 0.819$\pm$0.007 0.655$\pm$0.010 0.445$\pm$0.009 0.392$\pm$0.006
Imbalance-aware + Remix 0.918$\pm$0.006 0.819$\pm$0.007 0.693$\pm$0.009 0.588$\pm$0.007 0.431$\pm$0.008
+ Balanced-MixUp 0.910$\pm$0.007 0.800$\pm$0.008 0.685$\pm$0.012 0.575$\pm$0.010 0.426$\pm$0.007
Policy-based + RandAugment 0.880$\pm$0.008 0.770$\pm$0.010 0.692$\pm$0.007 0.552$\pm$0.008 0.410$\pm$0.008
Selective Mixup + GradMix (ours) 0.949$\pm$0.004 0.835$\pm$0.006 0.714$\pm$0.009 0.604$\pm$0.006 0.452$\pm$0.006
C.8 More Details on Analysis of Model Performance Results
Continuing from Sec. 6.1, we provide a more detailed analysis of the model performance results. Among the mixup-based methods, Manifold Mixup consistently shows good results. However, another mixup-based method, CutMix, shows the worst performance on CIFAR-10, CIFAR-100, and Tiny ImageNet as it significantly reduces accuracy on previous tasks. We believe that cutting images into patches in the limited previous task data results in significant information loss. The two imbalance-aware mixup methods, Remix and Balanced-MixUp, achieve slightly better results than Mixup as they address the data imbalance between previous and current tasks using disentangled mixing and balanced sampling techniques, respectively. The policy-based method, RandAugment, also shows underperforming results. For MNIST and FMNIST, elastic distortions across scale, position, and orientation are known to be effective, but other transformations in RandAugment like flipping and random cropping may have negative effects. For CIFAR-10, CIFAR-100, and Tiny ImageNet, the transformation technique in RandAugment that removes a region of an image is harmful to retaining the knowledge of the limited buffer data.
C.9 More Details on Sequential Performance Results
@@
Figure 12: Sequential accuracy results of GradMix and baselines on the five datasets as new tasks are learned.
C.10 More Details on the Effectiveness of GradMix on Large-Scale Datasets
Continuing from Sec. 6.1, we perform additional experiments to demonstrate the effectiveness of GradMix on large-scale datasets. We use the ImageNet-1K [18] dataset with 10 incremental tasks and provide the average accuracy results in Table 13. Using ER as a basis method, the experimental results show that GradMix continues to outperform other data augmentation baselines on the large-scale dataset. In general, GradMix is scalable by only computing the last layer gradients of a small exemplar set for each class.
Table 13: Average accuracy results on the ImageNet-1K dataset. We use ER as a basis method of class-incremental learning, and other experimental settings are the same as in Table 10.
@@ C.10 More Details on the Effectiveness of GradMix on Large-Scale Datasets
Experience Replay ER 0.326$\pm$0.012
Mixup-based + Mixup 0.313$\pm$0.013
+ Manifold Mixup 0.328$\pm$0.015
+ CutMix 0.281$\pm$0.016
Imbalance-aware + Remix 0.335$\pm$0.014
+ Balanced-MixUp 0.322$\pm$0.012
Policy-based + RandAugment 0.305$\pm$0.013
Selective Mixup + GradMix (ours) 0.356$\pm$0.010
C.11 More Details on Varying the Backbone Model
Continuing from Sec. 6.1, we perform additional experiments when using an initialized ViT [19] as the backbone model on the CIFAR-100 and Tiny ImageNet datasets. The average accuracy results are shown in Table 14, demonstrating the effectiveness of GradMix compared to the baselines when combined with ViT. However, the overall performance with ViT is lower than that with ResNet-18, which indicates that ViT may not be well-suited for continual learning settings with limited buffer data due to constraints in memory capacity, computational efficiency, and privacy. In addition, we believe that using a pre-trained model as the initial model would further improve overall performance, while GradMix would still be effectively adopted, as it relies on a gradient-based criterion adaptively derived from both data and the model.
Table 14: Average accuracy results on the CIFAR-100 and Tiny ImageNet datasets with ResNet-18 and ViT backbone models. We use ER as a basis method of class-incremental learning, and other experimental settings are the same as in Table 10.
@@ C.11 More Details on Varying the Backbone Model
Experience Replay ER 0.491$\pm$0.019 0.371$\pm$0.015 0.385$\pm$0.009 0.275$\pm$0.010
Mixup-based + Mixup 0.507$\pm$0.010 0.402$\pm$0.012 0.371$\pm$0.008 0.284$\pm$0.009
+ Manifold Mixup 0.521$\pm$0.008 0.410$\pm$0.011 0.386$\pm$0.009 0.298$\pm$0.011
+ CutMix 0.404$\pm$0.009 0.315$\pm$0.014 0.344$\pm$0.010 0.236$\pm$0.011
Imbalance-aware + Remix 0.528$\pm$0.007 0.426$\pm$0.010 0.391$\pm$0.008 0.305$\pm$0.009
+ Balanced-MixUp 0.516$\pm$0.009 0.396$\pm$0.013 0.379$\pm$0.008 0.281$\pm$0.010
Policy-based + RandAugment 0.484$\pm$0.007 0.353$\pm$0.012 0.368$\pm$0.007 0.263$\pm$0.008
Selective Mixup + GradMix (ours) 0.546$\pm$0.008 0.460$\pm$0.012 0.405$\pm$0.006 0.327$\pm$0.010
C.12 More Details on Selective Mixup Analysis
@@
Figure 13: Selective mixup results of GradMix on the last task of the five datasets.
C.13 More Details on Varying the Buffer Size
Continuing from Sec. 6.4, we provide more details on experimental results when varying the buffer size as shown in Table 15 and Table 16. As a result, GradMix outperforms other baselines regardless of the buffer size. For the results on the Tiny ImageNet dataset with a buffer size of 16 samples per class, we believe this buffer size is too small to train the model sufficiently, as none of the data augmentation methods improve accuracy compared to ER.
Table 15: Average accuracy results on the MNIST, FMNIST, and CIFAR-10 datasets with respect to the buffer size per class. We use ER as a basis method of class-incremental learning and compare GradMix with three types of data augmentation baselines. The buffer size is varied from 32 to 16 and 64 samples per class.
Method MNIST FMNIST CIFAR-10
$N$=16 $N$=64 $N$=16 $N$=64 $N$=16 $N$=64
ER 0.843$\pm$0.011 0.934$\pm$0.001 0.716$\pm$0.006 0.792$\pm$0.004 0.546$\pm$0.022 0.682$\pm$0.005
+ Mixup 0.807$\pm$0.016 0.921$\pm$0.002 0.733$\pm$0.006 0.811$\pm$0.005 0.569$\pm$0.022 0.718$\pm$0.010
+ Manifold Mixup 0.832$\pm$0.012 0.942$\pm$0.002 0.739$\pm$0.004 0.818$\pm$0.003 0.563$\pm$0.012 0.688$\pm$0.011
+ CutMix 0.798$\pm$0.014 0.933$\pm$0.004 0.743$\pm$0.011 0.830$\pm$0.003 0.475$\pm$0.009 0.602$\pm$0.016
+ Remix 0.848$\pm$0.014 0.944$\pm$0.004 0.751$\pm$0.006 0.825$\pm$0.005 0.565$\pm$0.017 0.711$\pm$0.011
+ Balanced-MixUp 0.823$\pm$0.013 0.937$\pm$0.002 0.724$\pm$0.005 0.818$\pm$0.005 0.560$\pm$0.028 0.696$\pm$0.006
+ RandAugment 0.689$\pm$0.016 0.833$\pm$0.003 0.600$\pm$0.008 0.696$\pm$0.005 0.584$\pm$0.008 0.703$\pm$0.010
+ GradMix (ours) 0.852$\pm$0.010 0.950$\pm$0.002 0.764$\pm$0.009 0.837$\pm$0.007 0.588$\pm$0.021 0.739$\pm$0.009
Table 16: Average accuracy results on the CIFAR-100 and Tiny ImageNet datasets with respect to the buffer size per class. Other experimental settings are the same as in Table 15.
     Method      CIFAR-100      Tiny ImageNet
     $N$=16      $N$=64      $N$=16      $N$=64
     ER      0.412$\pm$0.021      0.555$\pm$0.021      0.322$\pm$0.012      0.444$\pm$0.009
     + Mixup      0.413$\pm$0.014      0.587$\pm$0.016      0.281$\pm$0.009      0.446$\pm$0.011
     + Manifold Mixup      0.419$\pm$0.013      0.603$\pm$0.011      0.292$\pm$0.011      0.455$\pm$0.009
     + CutMix      0.347$\pm$0.007      0.505$\pm$0.006      0.289$\pm$0.009      0.414$\pm$0.010
     + Remix      0.423$\pm$0.010      0.607$\pm$0.012      0.299$\pm$0.009      0.460$\pm$0.009
     + Balanced-MixUp      0.426$\pm$0.016      0.597$\pm$0.016      0.286$\pm$0.010      0.449$\pm$0.012
     + RandAugment      0.415$\pm$0.013      0.560$\pm$0.009      0.304$\pm$0.007      0.450$\pm$0.010
     + GradMix (ours)      0.437$\pm$0.009      0.620$\pm$0.011      0.299$\pm$0.008      0.470$\pm$0.008
Appendix D Appendix – More Related Work
@@
Data augmentation is an effective method to enhance the generalization and robustness of models by generating additional training data. Traditional data augmentation techniques include image transformations such as random rotation, flipping, and cropping, which introduce variations to the input images [60, 71]. Recent works propose an automatic strategy to find an effective data augmentation policy [16, 17, 53]. However, these policy-based approaches do not consider continual learning scenarios, which may result in severe catastrophic forgetting as we show in our experiments.
@@
Another line of research is Mixup [75], which generates new synthetic data by linearly interpolating between pairs of features and their corresponding labels, and there are variant mixup-based methods [62, 73, 24]. In addition, imbalance-aware mixup methods have been proposed to improve model robustness on imbalanced datasets. A recent study [15] relaxes the mixing factor of labels to assign higher weights to the minority class while features are mixed in the same way as vanilla Mixup. Another study [21] selects sample pairs for mixup using both instance-based and class-based sampling in order to be robust for learning highly imbalanced medical image datasets. Furthermore, selective mixup is introduced by selecting sample pairs for mixup based on specific criteria rather than random, and it improves the model’s ability for domain adaptation [69] and out-of-distribution (OOD) generalization [72].
@@
A recent line of research applies existing data augmentation methods in class-incremental learning to further improve accuracy. The first work [50] simply combines vanilla Mixup to experience replay methods. Another work [6] employs the combination of existing data augmentation methods to enhance the diversity of samples in a buffer. In addition, a recent work [41] proposes adversarial perturbation technique that moves buffer data towards the current task data and utilizes Mixup within buffer data before adding perturbation for more diversity. Extending this line of work, [54] improves the sampling and mixing factor strategies in Mixup to address distribution discrepancies between the imbalanced training set and the balanced test set in continual learning, which is similar to existing imbalance-aware mixup methods [15, 21]. Another work [34] addresses class-incremental learning by using out-of-distribution detection methods that utilize image rotation as a data augmentation technique. Another line of research is using generative models to generate data for previous tasks [59, 22, 35], but this approach incurs additional costs for continuously training generative models across tasks. Also, there is a work [82] to achieve comparable performance with experience replay methods without storing buffer data by utilizing class augmentation and semantic augmentation techniques. In the setting of online continual learning, a recent study [78] repeatedly trains on buffer data for each incoming task data and uses random augmentation to alleviate overfitting caused by the repeated data. However, all these works simply use existing data augmentation techniques for further improvements and do not propose any novel data augmentation methods specifically designed for continual learning scenarios. In comparison, GradMix is a robust data augmentation method designed to minimize catastrophic forgetting in class-incremental learning. Although adversarial attack-based methods are studied for feature augmentation in class-incremental learning [37, 80], they are generally regarded as a distinct line of research compared to Mixup-based approaches in the literature [75, 62, 73, 24]. Since our work focuses specifically on Mixup techniques for class-incremental learning, we do not include adversarial attack-based methods in this study.

### v2

@@ B.5. More Details on Computational Complexity and Runtime Results of GradMix
Figure 12. Runtime results for solving the optimization problem in GradMix per epoch with respect to the number of tasks on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
In addition, we provide the average training time results over all tasks, both without and with GradMix, based on the experience replay method (ER) across all datasets as shown in Table 8. Note that other data augmentation baselines incur negligible computational overhead, resulting in training times similar to the method without GradMix. The computational overhead of GradMix depends on the last layer gradient computation, but our method computes gradients of only a small exemplar set for each class, not the whole training data. While GradMix incurs some additional computational cost relative to conventional training, we believe this is a worthwhile trade-off as it consistently leads to improved overall accuracy.
Table 8. Average training time results over all tasks, both without and with GradMix, based on the experience replay method (ER) on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
@@
B.6. More Details on Sensitivity of GradMix to the Mixup Parameter $\alpha$
Continuing from Sec. 6, we perform additional experiments to analyze the sensitivity of GradMix to the Mixup parameter $\alpha$ as shown in Table 9. As a result, the average accuracy of GradMix is not sensitive to $\alpha$. We believe that mixing the helpful pairs of classes that do not cause catastrophic forgetting might be beneficial to stabilize its performance regardless of $\alpha$. For Tiny ImageNet, a small value of $\alpha$ shows better accuracy results as in the Mixup paper (Zhang et al., 2018), and our selective mixup algorithm struggles to overcome the data characteristics.
Table 9. Average accuracy results of GradMix when varying the $\alpha$ value to 0.4, 1.0, and 2.0 on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
@@ B.6. More Details on Sensitivity of GradMix to the Mixup Parameter $\alpha$
[…]
B.7. More Details on Compatibility of GradMix to Advanced Experience Replay Methods
Continuing from Sec. 6.1, we perform additional experiments when using MIR (Aljundi et al., 2019a), GSS (Aljundi et al., 2019b), DER (Buzzega et al., 2020), and FOSTER (Wang et al., 2022) as basis experience replay methods for class-incremental learning, and the results are shown in Table 10–Table 13. As a result, GradMix still outperforms other data augmentation baselines when based on advanced experience replay methods. We believe that GradMix also can be extended to other experience replay methods including ASER (Shim et al., 2021) and SCR (Mai et al., 2021).
Table 10. Average accuracy results on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. We use MIR as a basis method of class-incremental learning and compare GradMix with three types of data augmentation baselines: (1) mixup-based methods: Mixup, Manifold Mixup, and CutMix; (2) imbalance-aware mixup methods: Remix and Balanced-MixUp; and (3) a policy-based method: RandAugment.
@@ B.7. More Details on Compatibility of GradMix to Advanced Experience Replay Methods
[…]
Table 11. Average accuracy results on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. We use GSS as a basis method of class-incremental learning, and other experimental settings are the same as in Table 10.
@@
[…]
Table 12. Average accuracy results on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. We use DER as a basis method of class-incremental learning, and other experimental settings are the same as in Table 10.
@@
[…]
Table 13. Average accuracy results on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. We use FOSTER as a basis method of class-incremental learning, and other experimental settings are the same as in Table 10.
Type Method MNIST FMNIST CIFAR-10 CIFAR-100 Tiny ImageNet
[…]
B.8. More Details on Sequential Performance Results
@@
Figure 13. Sequential accuracy results of GradMix and baselines on the five datasets as new tasks are learned.
B.9. More Details on the Effectiveness of GradMix on Large-Scale Datasets
Continuing from Sec. 6.1, we perform additional experiments to demonstrate the effectiveness of GradMix on large-scale datasets. We use the ImageNet-1K (Deng et al., 2009) dataset with 10 incremental tasks and provide the average accuracy results in Table 14. Using ER as a basis method, the experimental results show that GradMix continues to outperform other data augmentation baselines on the large-scale dataset. In general, GradMix is scalable by only computing the last layer gradients of a small exemplar set for each class.
Table 14. Average accuracy results on the ImageNet-1K dataset. We use ER as a basis method of class-incremental learning, and other experimental settings are the same as in Table 10.
@@ B.9. More Details on the Effectiveness of GradMix on Large-Scale Datasets
[…]
B.10. More Details on Varying the Backbone Model
Continuing from Sec. 6.1, we perform additional experiments when using an initialized ViT (Dosovitskiy et al., 2021) as the backbone model on the CIFAR-100 and Tiny ImageNet datasets. The average accuracy results are shown in Table 15, demonstrating the effectiveness of GradMix compared to the baselines when combined with ViT. However, the overall performance with ViT is lower than that with ResNet-18, which indicates that ViT may not be well-suited for continual learning settings with limited buffer data due to constraints in memory capacity, computational efficiency, and privacy. Nevertheless, transformer architectures specifically designed for continual learning, such as DyTox (Douillard et al., 2022), have demonstrated strong performance by dynamically expanding task-specific tokens, suggesting the potential of tailored transformer designs in overcoming the limitations observed with standard ViT backbones. In addition, we believe that using a pre-trained model as the initial model would further improve overall performance, while GradMix would still be effectively adopted, as it relies on a gradient-based criterion adaptively derived from both data and the model.
Table 15. Average accuracy results on the CIFAR-100 and Tiny ImageNet datasets with ResNet-18 and ViT backbone models. We use ER as a basis method of class-incremental learning, and other experimental settings are the same as in Table 10.
@@ B.10. More Details on Varying the Backbone Model
[…]
B.11. More Details on Selective Mixup Analysis
@@
Figure 14. Selective mixup results of GradMix on the last task of the five datasets.
Appendix C Appendix – More Related Work
@@
Data augmentation is an effective method to enhance the generalization and robustness of models by generating additional training data. Traditional data augmentation techniques include image transformations such as random rotation, flipping, and cropping, which introduce variations to the input images (Shorten and Khoshgoftaar, 2019; Yang et al., 2022). Recent works propose an automatic strategy to find an effective data augmentation policy (Cubuk et al., 2019, 2020; Müller and Hutter, 2021). However, these policy-based approaches do not consider continual learning scenarios, which may result in severe catastrophic forgetting as we show in our experiments.
@@
Another line of research is Mixup (Zhang et al., 2018), which generates new synthetic data by linearly interpolating between pairs of features and their corresponding labels, and there are variant mixup-based methods (Verma et al., 2019; Yun et al., 2019; Harris et al., 2021). In addition, imbalance-aware mixup methods have been proposed to improve model robustness on imbalanced datasets. A recent study (Chou et al., 2020) relaxes the mixing factor of labels to assign higher weights to the minority class while features are mixed in the same way as vanilla Mixup. Another study (Galdran et al., 2021) selects sample pairs for mixup using both instance-based and class-based sampling in order to be robust for learning highly imbalanced medical image datasets. Furthermore, selective mixup is introduced by selecting sample pairs for mixup based on specific criteria rather than random, and it improves the model’s ability for domain adaptation (Xu et al., 2020) and out-of-distribution (OOD) generalization (Yao et al., 2022).
@@
A recent line of research applies existing data augmentation methods in class-incremental learning to further improve accuracy. The first work (Mi et al., 2020) simply combines vanilla Mixup to experience replay methods. Another work (Bang et al., 2021) employs the combination of existing data augmentation methods to enhance the diversity of samples in a buffer. In addition, a recent work (Kumari et al., 2022) proposes adversarial perturbation technique that moves buffer data towards the current task data and utilizes Mixup within buffer data before adding perturbation for more diversity. Extending this line of work, (Qiang et al., 2023) improves the sampling and mixing factor strategies in Mixup to address distribution discrepancies between the imbalanced training set and the balanced test set in continual learning, which is similar to existing imbalance-aware mixup methods (Chou et al., 2020; Galdran et al., 2021). Another work (Kim et al., 2022) addresses class-incremental learning by using out-of-distribution detection methods that utilize image rotation as a data augmentation technique. Another line of research is using generative models to generate data for previous tasks (Shin et al., 2017; Gao and Liu, 2023; Kim et al., 2024a), but this approach incurs additional costs for continuously training generative models across tasks. Also, there is a work (Zhu et al., 2021) to achieve comparable performance with experience replay methods without storing buffer data by utilizing class augmentation and semantic augmentation techniques. In the setting of online continual learning, a recent study (Zhang et al., 2022b) repeatedly trains on buffer data for each incoming task data and uses random augmentation to alleviate overfitting caused by the repeated data. However, all these works simply use existing data augmentation techniques for further improvements and do not propose any novel data augmentation methods specifically designed for continual learning scenarios. In comparison, GradMix is a robust data augmentation method designed to minimize catastrophic forgetting in class-incremental learning. Although adversarial attack-based methods are studied for feature augmentation in class-incremental learning (Kim et al., 2024c; Zheng et al., 2024), they are generally regarded as a distinct line of research compared to Mixup-based approaches in the literature (Zhang et al., 2018; Verma et al., 2019; Yun et al., 2019; Harris et al., 2021). Since our work focuses specifically on Mixup techniques for class-incremental learning, we do not include adversarial attack-based methods in this study.
Appendix D Appendix – Limitations
The computational overhead of GradMix inevitably grows as the number of classes increases over time in class-incremental learning. Despite this limitation, we believe this is a worthwhile trade-off as GradMix consistently achieves improved overall accuracy. However, GradMix may not be effective in tasks where mixing samples is impractical, such as semantic segmentation or tasks involving discrete symbolic features, where interpolating either the feature or the label may lead to semantically inconsistent samples.